OpenStudy (anonymous):

Please help! Given f(x) = x/x+3 find f(f(x)). I tried to solve it and I got 7/12? Is this correct? If not please explain it to me step by step.

OpenStudy (amriju):

miss andrea..how come you get avlue when none is provided?

OpenStudy (anonymous):

I got you, hold up

OpenStudy (amriju):

* a value

OpenStudy (anonymous):

oops I have to find f(f(3))

OpenStudy (anonymous):

what is \(f(3)\) ?

OpenStudy (amriju):

step 1: find f(3) step 2: remember the value of f(3), let it be a. step 3: find f(a).

OpenStudy (amriju):

the answer i guess is 1/7..

OpenStudy (anonymous):

aren't I supposed to plug in 3 into every x?

OpenStudy (anonymous):

f(f(x)) means that x values in f(x) is equal to the f(x) equation.

OpenStudy (anonymous):

I have no idea what that means, u just made this a million times more confusing

OpenStudy (anonymous):

You see that x is divided by x, so.......what happens if you switch the x with the f(x) equation, theyy.........

OpenStudy (amriju):

no m'lady..u are supposed to put in the inner function, then get a value of the inner function..and put the value in the outer function

OpenStudy (anonymous):

If you divide same thing by the same thing they........

OpenStudy (amriju):

*supposed to put x in the inner function...

OpenStudy (amriju):

so...basically u find f(3)..and then put its value in the outer function, that is also f..thats why f(f(x))

OpenStudy (anonymous):

i have no idea what that means

OpenStudy (anonymous):

@amriju would the answer be 6/7 ?

OpenStudy (anonymous):

@Flashy32 would the answer be 6/7?

OpenStudy (amriju):

f(3)=3/6=.5 f(.5)=.5/3.5=1/7

OpenStudy (anonymous):

@amriju thanks, i just saw my mistake !

OpenStudy (anonymous):

@amriju btw could u help me with one more?

OpenStudy (amriju):

sure